OpenStudy (anonymous):

can someone help me understant multiplying and dividing rational expressions ? i dont really understand it, and need too.

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{\frac{a}{b} \times \frac{c}{d}={ac \over bd}}}\]\[\Huge{\color{green}{\frac{a}{b} \div \frac{c}{d}=\frac{a}{b} \times \frac{c}{d}=\frac{ac}{bd}}}\]